# Semilinear diffusion equations on infinite graphs: the dissipative and Lipschitz cases

### v1

@@ 6.3.1 Upper estimates for the $u_{k}^{(n)}$’s when $q\in(0,1)$
The sequence $\{\theta_{k}\}_{k=1}^{N}$ defined in (6.4) is well-defined and satisfies the relation $\theta_{k}=\psi_{q,\lambda_{k}}^{-1}(\theta_{k-1})$. In particular, the implicit Euler system (6.4) admits a solution and the corresponding interpolant $\theta_{\epsilon}$ is an $\epsilon$-approximation to problem (6.5). Using the fact that $\operatorname{sgn}(F(\theta)-F(\eta))=\operatorname{sgn}(\theta-\eta)$ by monotonicity, it is not difficult to check that $F$ is an $m$-accretive operator on $\mathbb{R}$. Thus, by [3, Theorem 4.1], $\theta_{\epsilon}$ converges to the unique mild solution to problem (6.5). Thus, uniqueness of mild solutions for $m$-accretive $F$, yields $|\theta_{\epsilon}(t)-\theta(t)|\to 0$ as $\epsilon\to 0$, for every fixed tt.∎
@@
Since $u_{0}\geq 0$, by a recursive argument and Theorem 4.3-$(a)$, which obviously holds for finite graphs, $u_{k}^{(n)}\geq 0$. Let $x_{k}\in X_{n}$ be such that $u_{k}^{(n)}(x_{k})=\|u_{k}^{(n)}\|_{\infty}$. Since $G_{\mathrm{dir},n}$ is finite, such a node exists and, since $u_{k}^{(n}$ attains its maximum at $x_{k}$ $\Delta_{\mathrm{dir},n}u_{k}^{(n)}(x_{k})\geq 0$.
@@ 6.3.2 Positivity of solutions in the case $q\in(1,+\infty)$
For every fixed $n$, consider on the finite graph $G_{\operatorname{dir},n}$ the “mild” solution $u^{(n)}$ to which the sequence $\{u_{\epsilon}^{(n)}\}_{\epsilon}$ converges in $\ell^{p}(X,\mu)$ for all $t\in[0,T]$. For finite graphs, mild, strong and classic solutions are equivalent. Since $q\in(1,+\infty)$, from Theorem [10, Theorems 4.7-4.8], $u^{(n)}>0$ in $X_{n}$ and therefore there exists $\epsilon=\epsilon(n)$ such that for every $0<\epsilon\leq\epsilon(n)$ it holds
@@
Fix $x\in X$ and $t\in[0,T]$. Let $X_{n_{0}}$ be the first element of the exhaustion $\{X_{n}\}_{n}$ such that $x\in X_{n_{0}}$. By the inequality above, recalling that the sequence $\{u_{\epsilon}\}_{\epsilon}$ converges to the mild solution $u$ in $\ell^{p}(X,\mu)$ for all $t\in[0,T]$, we get the following pointwise estimate:
@@
We observe that, under the assumption of Theorem 3.6, the proof of extinction in finite time when $q\in(0,1)$ can be adapted to the case of a sign changing initial datum $u_{0}$ in the following way. Consider the sequences $\{u_{k}^{(n),+}\}_{k}$ and $\{v_{k}^{(n),-}\}_{k}$ defined recursively by (6.6) with starting value, respectively $u_{0}^{+}\geq 0$ and $u_{0}^{-}\geq 0$ (see (4.17)). Namely,
@@
Applying Lemma 6.7, we infer that $u_{k}^{(n),+}\leq\theta_{k}$ and $v_{k}^{(n),-}\leq\theta_{k}$. Furthermore, since $\bm{\pi}_{n}u_{0}^{+}\geq\bm{\pi}_{n}u_{0}$ and $\bm{\pi}_{n}u_{0}^{-}\geq\bm{\pi}_{n}u_{0}$, from Corollary 4.4 applied recursively, we also get that $u_{k}^{(n)}\leq u_{k}^{(n),+}$ and $u_{k}^{(n)}\leq v_{k}^{(n),-}$. Finally, we set $u_{k-1}^{(n),-}=-v_{k-1}^{(n),-}$ and we conclude that
@@
Hence the conclusion follows as in the proof of Theorem 3.6.

### v3

@@ 6.3.1 Upper estimates for the $u_{k}^{(n)}$’s when $q\in(0,1)$
The sequence $\{\theta_{k}\}_{k=1}^{N}$ defined in (6.4) is well-defined and satisfies the relation $\theta_{k}=\psi_{q,\lambda_{k}}^{-1}(\theta_{k-1})$. In particular, the implicit Euler system (6.4) admits a solution and the corresponding interpolant $\theta_{\epsilon}$ is an $\epsilon$-approximation to problem (6.5). Using the fact that $\operatorname{sgn}(F(\theta)-F(\eta))=\operatorname{sgn}(\theta-\eta)$ by monotonicity, it is not difficult to check that $F$ is an $m$-accretive operator on $\mathbb{R}$. Thus, by [3, Theorem 4.1], $\theta_{\epsilon}$ converges to the unique mild solution to problem (6.5). Thus, uniqueness of mild solutions for $m$-accretive $F$, yields $|\theta_{\epsilon}(t)-\theta(t)|\to 0$ as $\epsilon\to 0$, for every fixed $t$. ∎
@@
Since $u_{0}\geq 0$, by a recursive argument and Theorem 4.3-$(a)$, which obviously holds for finite graphs, $u_{k}^{(n)}\geq 0$. Let $x_{k}\in X_{n}$ be such that $u_{k}^{(n)}(x_{k})=\|u_{k}^{(n)}\|_{\infty}$. Since $G_{\mathrm{dir},n}$ is finite, such a node exists and, since $u_{k}^{(n)}$ attains its maximum at $x_{k}$ $\Delta_{\mathrm{dir},n}u_{k}^{(n)}(x_{k})\geq 0$.
@@ 6.3.2 Positivity of solutions in the case $q\in(1,+\infty)$
For every fixed $n$, consider on the finite graph $G_{\operatorname{dir},n}$ the “mild” solution $u^{(n)}$ to which the sequence $\{u_{\epsilon}^{(n)}\}_{\epsilon}$ converges in $\ell^{p}(X,\mu)$ for all $t\in[0,T]$. For finite graphs, mild, strong and classic solutions are equivalent. Since $q\in(1,+\infty)$, from [10, Theorems 4.7-4.8], $u^{(n)}>0$ in $X_{n}$ and therefore there exists $\epsilon=\epsilon(n)$ such that for every $0<\epsilon\leq\epsilon(n)$ it holds
@@
Fix $x\in X$ and $t\in[0,T]$. Let $X_{n_{0}}$ be the first element of the exhaustion $\{X_{n}\}_{n}$ such that $x\in X_{n_{0}}$. By the inequality above, and since the convergence of the sequence $\{u_{\epsilon}\}_{\epsilon}$ to the mild solution $u$ in $\ell^{p}(X,\mu)$ implies the pointwise convergence for every fixed $t\in[0,T]$ and $x\in X$ , we get the following pointwise estimate:
@@
We observe that, under the assumption of Theorem 3.6, the proof of extinction in finite time when $q\in(0,1)$ can be adapted to the case of a sign changing initial datum $u_{0}\in\ell^{p}(X,\mu)\cap\ell^{\infty}(X,\mu)$ in the following way. Consider the sequences $\{u_{k}^{(n),+}\}_{k}$ and $\{v_{k}^{(n),-}\}_{k}$ defined recursively by (6.6) with starting value, respectively $u_{0}^{+}\geq 0$ and $u_{0}^{-}\geq 0$ (see (4.18)). Namely,
@@
Applying Lemma 6.6, we infer that $u_{k}^{(n),+}\leq\theta_{k}$ and $v_{k}^{(n),-}\leq\theta_{k}$. Furthermore, since $\bm{\pi}_{n}u_{0}^{+}\geq\bm{\pi}_{n}u_{0}$ and $\bm{\pi}_{n}u_{0}^{-}\geq-\bm{\pi}_{n}u_{0}$, from Corollary 4.4 applied recursively, we also get that $u_{k}^{(n)}\leq u_{k}^{(n),+}$ and $u_{k}^{(n)}\geq-v_{k}^{(n),-}$. Finally, we set $u_{k}^{(n),-}=-v_{k}^{(n),-}$ and we conclude that $u_{k}^{(n),-}\leq u_{k}^{(n)}\leq u_{k}^{(n),+}$ for $k=0,\ldots,N$, and, for the corresponding piecewise-constant interpolants,
@@
Hence the conclusion follows as in the proof of Theorem 6.2-(1).